What is the area of a right triangle that has sides that are 3m, 4m and 5m?
Answer by Alan3354(21583)

What is the area of a right triangle that has sides that are 3m, 4m and 5m?
------
Area = bh/2
A = 3*4/2 = 6 sq meters